1, Explain the role of a salt bridge in a voltaic cell

2, Assign oxidation numbers to the following compounds:

a. KClO3 b. NO3 c. CH3 CHO

3, Balance the following half-reaction in an acidic solution:

Cr2 O72- → Cr3+

1
Expert's answer
2021-01-11T03:47:17-0500
  1. Salt bridges complete the electrical circuit. They also allow the movement of ions from one solution to another without mixing of the two solutions and maintain electrical neutrality.
  2. KClO3: oxidation number of K is +1, oxidation number of Cl is +5, oxidation number of O is -2, the oxidation number of the compound (the sum) is 1+5-2*3=0. NO3: oxidation number of N is +5, oxidation number of O is -2, the oxidation number of the compound (the sum) is 5-2*3=-1. C1H13C2H2O: oxidation number of C1 is -3, oxidation number of H1 is +1, oxidation number of C2 is +1, the oxidation number of H2 is +1, oxidation number of O is -2, the oxidation number of the compound (the sum) is -3+1*3+1+1-2=0.
  3. Cr2O72– + 7H+ + 4e → 2Cr3+ + 3.5H2O